Suppose two dice are rolled. Find the probabilities of the following events.
a) the maximum of the two numbers rolled is less than or equal to 2;
b) the maximum of the two numbers rolled is less than or equal to 3;
c) the maximum of the two numbers rolled is exactly equal to 3.
d) Repeatb)andc)forxinsteadcf3,foreachxfrom1to6.
e) Denote by P(x) the probability that the maximum number is exactly x. What should P(I) +P(2) +P(3) +P(4) +P(5) +P(6) equal? Check this for your answers to d).
Answer #1

a)

X be highest of two values
outcome X
(1,1) 1
(1,2)(2,1)(2,2) 2

required probability = 4/36 = 1/9

b)

X be highest of two values
outcome X
(1,1) 1
(1,2)(2,1)(2,2) 2
(1,3)(3,1)(2,3)(3,2)(3,3) 3

total outcome=9

required probability =9/36 = 1/4

c)

outcome X
(1,3)(3,1)(2,3)(3,2)(3,3) 3

required probability = 5/36

d)

for b)

let two numbers rolled is less than or equal to X

X P(X)
≤1 1/36
≤2 4/36
≤3 9/36
≤4 16/36
≤5 25/36
≤6 36/36

for c)

X be highest of two values
outcome X P(X)
(1,1) 1 1/36
(1,2)(2,1)(2,2) 2 3/36
(1,3)(3,1)(2,3)(3,2)(3,3) 3 5/36
(1,4)(4,1)(2,4)(4,2)(3,4)(4,3)(4,4) 4 7/36
(1,5)(5,1)(2,5)(5,2)(3,5)(5,3)(4,5)(5,4)(5,5) 5 9/36
(1,6)(6,1)(2,6)(6,2)(3,6)(6,3)(4,6)(6,4)(5,6)(6,5)(6,6) 6 11/36

e)

X P(X)
1 1/36
2 3/36
3 5/36
4 7/36
5 9/36
6 11/36

P(I) +P(2) +P(3) +P(4) +P(5) +P(6) = 36/36 = 1